# Finite-time convergence to an $\epsilon$-efficient Nash equilibrium in potential games

### v2

@@ Abstract
This paper investigates the convergence time of log-linear learning to an $\epsilon$-efficient Nash equilibrium (NE) in potential games. In such games, an efficient NE is defined as the maximizer of the potential function. Existing results are limited to potential games with stringent structural assumptions and entail exponential convergence times in $1/\epsilon$. Unaddressed so far, we tackle general potential games and prove the first finite-time convergence to an $\epsilon$-efficient NE. In particular, by using a problem-dependent analysis, our bound depends polynomially on $1/\epsilon$. Furthermore, we provide two extensions of our convergence result: first, we show that a variant of log-linear learning that requires a factor $A$ less feedback on the utility per round enjoys a similar convergence time; second, we demonstrate the robustness of our convergence guarantee if log-linear learning is subject to small perturbations such as alterations in the learning rule or noise-corrupted utilities.
@@ 1 Introduction
Interactions of multiple agents are at the heart of many applications in transportation networks, auctions, telecommunication networks, and multi-robot systems. In game theory, the Nash equilibrium is a popular solution concept to describe outcomes of a multi-agent system [27]. Fundamental considerations in game theory are whether an NE exists, if strategic players can learn it, and if so at which speed they can learn it. Furthermore, for practical purposes it is important to understand which NE is learned. This is particularly pertinent in games that admit a social welfare function, as it enables the definition of an efficient NE as an NE that maximizes the social welfare. The social welfare is typically an aggregate measure of individual utilities such as their sum or a measure of fairness. In distributed control of engineering systems, for example, studying efficient NEs is important, as the aim is to optimize some global objective function in a distributed manner.
The class of potential games [23] lends itself to studying efficient NEs since every joint action maximizing the potential corresponds to an NE. This property follows from the fact that in potential games, the difference in a player’s utility generated by a unilateral change of her action equals the difference in potential. Consequently, if the social welfare function is aligned with the potential function, meaning that an increase in social welfare is associated with an increase in potential [29], then any joint action maximizing the potential is an efficient NE [18]. In identical interest games, for example, maximizing the aggregated utility is achieved by maximizing the potential function which is trivially given by the common utility function. In the example of coverage problems, the goal is to assign players to locations to achieve maximal coverage. When each player’s utility is designed as the marginal contribution then this goal is achieved by maximizing the resulting potential function [19].
In this paper, we investigate the speed at which an efficient NE can be reached in potential games.
@@ 1.1 Related work
In potential games, several learning rules exist such as iterative best-response dynamics [31, 4, 9], no-regret algorithms [13, 30, 11], and fictitious play [22, 23] for which asymptotic convergence to an NE is guaranteed. However, only log-linear learning [5, 33] and variants thereof [1, 20, 17] are known to converge to an efficient Nash equilibrium. In log-linear learning, the players asynchronously choose an action with a probability proportional to its exponentiated utility. To increase its applicability, existing works also studied log-linear learning under more realistic structural assumptions. For example, Leslie and Marden [14] proves that log-linear learning also handles the practical setting where the observed utilities are corrupted by noise. Furthermore, Arslan et al. [1] proposes binary log-linear learning, a slight modification of log-linear learning that only requires two points of feedback per round. Although the above works established asymptotic convergence of log-linear learning to an efficient NE, a finite-time analysis is missing for general potential games.
Few past works provide finite-time guarantees for log-linear learning to an $\epsilon$-efficient NE, an action profile whose potential is $\epsilon$-close to its maximum value. For instance, for atomic routing games with polynomial costs of degree at most $p$, Asadpour and Saberi [2] proves a convergence time exponential in $p$ and $1/\epsilon$ and polynomial in $N$, where $N$ denotes the number of players. Moreover, Montanari and Saberi [24, 25] study games with graph structures between players and prove an exponential convergence time in $N$ and $1/\epsilon$ in the worst case. Finally, in potential games with interchangeable players and a Lipschitz-continuous potential function, Shah and Shin [32] shows a convergence time exponential in $A$ and $1/\epsilon$ and linear in $N$, where $A$ is the number of actions per player.
For the class of games we address in this paper, namely general potential games, an exponential dependence on $N$ in the convergence time is unavoidable. In fact, finding an $\epsilon$-efficient NE is equivalent to finding an $\epsilon$-optimizer of the potential function, which is NP-complete [8]. However, there is no hardness result that justifies the exponential dependence on $1/\epsilon$ present in previous works. In this paper, we focus on deriving the first finite-time convergence guarantees of log-linear learning that hold for general potential games; are polynomial in $1/\epsilon$; and remain valid under relaxed structural assumptions such as limited access to feedback and noisy utilities.
@@ 1.2 Contributions
  • We prove a convergence time of $\tilde{\mathcal{O}}((A^{N}/\epsilon)^{\frac{1}{\max\{\epsilon,\Delta\}}})$ to an $\epsilon$-efficient NE (Theorem 1), where $\Delta$ is a problem-dependent constant. If in addition, the players are interchangeable, then an $\epsilon$-NE is reached in $\tilde{\mathcal{O}}((\frac{N^{A}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$ which in contrast to general potential games is polynomial in $N$ as well (Corollary 4).
  • We consider two variants, binary log-linear learning (Theorem 2) and perturbed log-linear learning (Theorem 3) motivated by limited feedback and noise corrupted utilities, respectively. For these variants we show convergence guarantees that are polynomial in $1/\epsilon$.
On the technical side, in their convergence analysis, most past works leveraged that log-linear learning induces a Markov chain. To obtain our novel finite-time results we build on this connection and develop new Markov chain results that can be summarized as follows:
  • We derive improved mixing time bounds for a given class of Markov chains based on log-Sobolev inequalities (Lemma 1). In particular, this broad class of Markov chains includes those induced by log-linear learning and binary-log linear learning.
  • We derive a tight Lipschitz constant for the known result regarding the Lipschitz-continuity of stationary distributions of Markov chains as a function of their transition matrix (Lemma 4). We leverage this result to study the convergence of learning rules such as perturbed log-linear learning for which the explicit stationary distributions are unknown (Theorem 3).
@@ 2.1 Problem setup
In this paper, we assume that an initial action profile is drawn from a distribution $\mu^{0}$. The potential game is then repeated over multiple rounds. We assume a turn-based setting, where at round $t$, a player $i\in[N]$ is uniformly selected to update her action while the other players remain with their previous action, i.e., $a_{-i}^{t}=a_{-i}^{t-1}$.
A common assumption in game theory is that the players are rational, i.e., that they seek to maximize their utility. Under this assumption, a natural solution concept is the pure Nash equilibrium [28].
Definition 1 (Nash equilibrium).
@@
In a pure NE, no player can improve her utility by unilaterally changing her action. Hereafter, we simply refer to a pure Nash equilibrium as a Nash equilibrium. A stronger solution concept specific to potential games is the efficient Nash equilibrium. An efficient Nash equilibrium is an action profile $a^{*}$ which maximizes the potential function, i.e., $a^{*}\in\operatorname*{arg\,max}_{a\in A^{N}}\Phi(a)$. An appealing property of finite potential games is that a potential maximizer exists, and thus an efficient NE exists [23].
Practically, each player selects her actions based on some learning rule. In this work, we are interested in learning rules that converge, in expectation, to an $\epsilon$-efficient NE, i.e., an action profile maximizing the potential up to an additive constant:
@@
If all players apply a learning rule that relies exclusively on the utility given the most recent action profile of the other players then the considered game induces a Markov chain over the state space $\mathcal{A}^{N}$. In particular, the state of the Markov chain at time $t$ corresponds to the action profile at round $t$ and the learning rule of each player specifies the transition dynamics. This connection between the learning dynamics in potential games and Markov chains has been exploited in previous works [5, 33, 17, 32] and is crucial for our subsequent convergence analysis. We provide the necessary background on Markov chains in the next section.
@@ 2.2 Background on Markov chains
We briefly review concepts and properties of Markov chains used throughout this paper. Consider a time-homogeneous Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ over the state space $\mathcal{A}^{N}$ with a transition matrix $P\in\mathbb{R}^{{A^{N}}\times A^{N}}$. The ergodic theorem [15] states that if a Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ is irreducible and aperiodic, then it has a unique stationary distribution $\mu$, and from any initial distribution $\mu^{0}$ the distribution $\mu^{t}=\mu^{0}P^{t}$ converges to $\mu$. The convergence time to the stationary distribution is quantified by the mixing time:
@@
where the total variation distance is defined as $\|\mu^{t}-\mu\|_{TV}:=\frac{1}{2}\sum_{a\in\mathcal{A}^{N}}|\mu^{t}(a)-\mu(a)|$. Based on a remark in [10, Section 3] we now derive a bound on the mixing time of Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$.
Lemma 1.
@@
We briefly discuss the mixing time bound above and provide a proof and a thorough discussion in Appendix A. While classical approaches commonly bound the mixing time by the spectral gap defined in Equation (11) in Appendix A, bounds using log-Sobolev constants are often significantly tighter. Indeed, the mixing time upper bound using the log-Sobolev constant grows as $\mathcal{O}(\log\log(1/\mu_{\min}))$ whereas bounds using the spectral gap grow as $\mathcal{O}(\log(1/\sqrt{\mu_{\min}}))$. However, unlike the spectral gap deriving log-Sobolev constants can be extremely difficult. So constants have not been well-explored. In the next section, we derive a novel bound on the log-Sobolev constant of a class of Markov chains.
3 Log-linear learning
In this section, we review the well-established log-linear learning rule [5] and state our main theoretical result on the convergence time of log-linear learning to an $\epsilon$-efficient Nash equilibrium. Our convergence analysis relies on the mixing time bound from Lemma 1 of the previous section.
@@ 3.1 Algorithm and background
where parameter $\beta$ measures a player’s rationality: for large $\beta$ player $i$ is more likely to select a best response $a_{i}^{t}\in\operatorname*{arg\,max}_{a_{i}\in\mathcal{A}}U_{i}(a_{i},a_{-i}^{t-1})$; and for $\beta=0$ player $i$ samples $a_{i}^{t}$ uniformly. Moreover, the strategy $p_{i}^{t}$ is myopic as it depends only on the other players’ last actions $a_{-i}^{t-1}$.
Log-linear learning induces an irreducible and aperiodic Markov chain $\{X_{t}\}_{t\in\mathbb{Z}_{+}}$ with a time-reversible transition matrix $P\in\mathbb{R}^{\mathcal{A}\times\mathcal{A}}$ [17] given by:
@@
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. The stationary distribution $\mu\in\Delta(\mathcal{A}^{N})$ of log-linear learning is given by [5]:
@@
For sufficiently large $\beta$, it was shown that log-linear learning converges asymptotically to a potential maximizer and thus to an efficient Nash equilibrium [5, 33, 17]. Except for a few works [25, 2, 32], none of the previous works, however, have finite-time convergence guarantees to such an efficient Nash equilibrium, and [25, 2, 32] have additional assumptions on the potential game. Thus, in the following section, we establish a bound on the convergence time of log-linear learning in general potential games.
3.2 Convergence time of log-linear learning
@@
In other words, after $t=\tilde{\Omega}(N^{2}A^{5}(\frac{A^{N}}{\epsilon})^{1/\max\{\epsilon,\Delta\}})$ rounds of log-linear learning with $\beta=\Omega(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{A^{N}}{\epsilon})$ the expected value of the potential of the joint action at time $t$ is $\epsilon$-optimal.
Theorem 1 provides the first finite-time convergence rates to an $\epsilon$-efficient NE in general potential games. In order to be $\epsilon$-close to an efficient NE $\beta$ must scale as $\Omega(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{A^{N}}{\epsilon})$. Furthermore, since the convergence time scales as $e^{4\beta}$ it grows polynomially in $\mathcal{A}$ and $1/\epsilon$ and exponentially in $N$. The exponential dependence on $N$ is unavoidable without further assumptions on the game. However, by deriving problem-dependent bounds we are the first to avoid the exponential dependence in $1/\epsilon$, see Table 1. Note that the case $\Delta=0$ is trivial as all action profiles are efficient NEs.
@@
Works Game setting Assumptions Convergence time
[2] Routing game Cost functions of degree $\tilde{\Omega}(e^{\frac{N}{\epsilon}})$
with $K$ vertices at most $p$
[32] Potential game with $\lambda$-Lipschitz continuous $\tilde{\Omega}(N(\frac{A\lambda}{\epsilon})^{\frac{A}{\epsilon}})$
interchangeable players potential function
Corollary 4 Potential game with $A\geq 4$ $\tilde{\Omega}(N(\frac{N^{A}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$
interchangeable players
Theorem 1 Potential game $A\geq 4$ $\tilde{\Omega}(N^{2}A^{5}(\frac{A^{N}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$
Proof outline.
Here, we provide an outline of the proof, the full proof is deferred to Appendix B.1. The argument begins with the following decomposition:
@@
Second term: We can control the mixing time of log-linear learning using the bound of Lemma 1:
@@
Lemma 3.
Consider a Markov chain over state space $\mathcal{A}^{N}$ with $A\geq 4$. Assume that there exists $p_{\min},p_{\max}\in(0,1]$, such that the corresponding transition matrix $P\in\mathbb{R}^{A^{N}\times A^{N}}$ satisfies:
@@
for $t\geq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\left(\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\right)$. This concludes the proof of the theorem. ∎
In the following, we additionally assume that the potential game is symmetric, i.e., players are interchangeable. [^3] A definition of a symmetric game is given in Appendix B.4, see Definition 3. In this setting, we derive a bound on the convergence time of log-linear learning that depends polynomially not only on $1/\epsilon$ but also on $N$.
Corollary 1 (Sketch).
Consider a symmetric potential game. For any $\epsilon\in(0,1)$ and initial distribution $\mu^{0}$, assume that players adhere to modified log-linear learning [32] with $\beta=\Omega(\frac{1}{\max\{\epsilon,\Delta\}}\log(\frac{N^{A}}{\epsilon}))$. Then, $\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon$ for $t=\tilde{\Omega}(N(\frac{N^{A}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$.
We provide a full statement of the corollary and its proof in Appendix B.4. For symmetric potential games with a $\lambda$-Lipschitz-continuous potential function, Shah and Shin [32] proves a convergence time of $\tilde{\Omega}(N(\frac{A\lambda}{\epsilon})^{\frac{A}{\epsilon}})$. In comparison, relaxing the Lipschitzness assumption comes at a small cost of polynomial dependence on $N$ rather than a linear one. However, our result greatly improves the dependence on $\epsilon$ to a polynomial one.
4 Relaxed structural assumption
In this section, we investigate how our convergence time results are affected when structural assumptions are relaxed. In particular, we analyze binary log-linear learning which handles reduced feedback, then we study perturbations of log-linear learning such as noisy utility observations.

### v3

@@ Abstract
This paper investigates the convergence time of log-linear learning to an $\epsilon$-efficient Nash equilibrium (NE) in potential games. In such games, an efficient NE is defined as the maximizer of the potential function. Previous literature provides asymptotic convergence rates to efficient Nash equilibria, and existing finite-time rates are limited to potential games with further assumptions such as the interchangeability of players. In this paper, we prove the first finite-time convergence to an $\epsilon$-efficient NE in general potential games. Our bounds depend polynomially on $1/\epsilon$, an improvement over previous bounds that are exponential in $1/\epsilon$ and only hold for subclasses of potential games. We then strengthen our convergence result in two directions: first, we show that a variant of log-linear learning that requires a factor $A$ less feedback on the utility per round enjoys a similar convergence time; second, we demonstrate the robustness of our convergence guarantee if log-linear learning is subject to small perturbations such as alterations in the learning rule or noise-corrupted utilities.
keywords:
Game theory, potential games, log-linear learning, efficient Nash equilibrium.
@@ 1 Introduction
Interactions of multiple agents are at the heart of many applications in transportation networks, auctions, telecommunication networks, and multi-robot systems. In game theory, the Nash equilibrium is a popular solution concept to describe outcomes of a multi-agent system (Nash, 1951). Fundamental considerations in game theory are whether an NE exists, if strategic players can learn it, and if so at which speed they can learn it. Furthermore, for practical purposes it is important to understand which NE is learned. This is particularly pertinent in games that admit a social welfare function, as it enables the definition of an efficient NE as an NE that maximizes social welfare. The social welfare is typically an aggregate measure of individual utilities such as their sum or a measure of fairness. In distributed control of engineering systems, for example, studying efficient NEs is important, as the aim is to optimize some global objective function in a distributed manner.
The class of potential games (Monderer and Shapley, 1996b) lends itself to studying efficient NEs since every joint action maximizing the potential corresponds to an NE. This property follows from the fact that in potential games, the difference in a player’s utility generated by a unilateral change of her action equals the difference in potential. Consequently, if the social welfare is aligned with the potential function, meaning that an increase in social welfare is associated with an increase in potential (Paccagnan et al., 2022), then any joint action maximizing the potential is an efficient NE (Marden and Shamma, 2015). In identical interest games, for example, maximizing the aggregated utility is achieved by maximizing the potential function which is trivially given by the common utility function. In the example of coverage problems, the goal is to assign players to locations to achieve maximal coverage. When each player’s utility is designed as the marginal contribution then this goal is achieved by maximizing the resulting potential function (Marden and Wierman, 2013).
In this paper, we study the speed of convergence to approximate efficient NE in potential games.
@@ 1.1 Related work
In potential games, many learning rules are proven to converge to an NE, such as iterative best-response dynamics (Rosenthal, 1973; Awerbuch et al., 2008; Chien and Sinclair, 2011), no-regret algorithms (Krichene et al., 2015; Palaiopanos et al., 2017; Heliou et al., 2017), and fictitious play (Monderer and Shapley, 1996a, b). However, only log-linear learning (Blume, 1993; Young, 1993) is known to converge to an efficient Nash equilibrium. [^1] In log-linear learning, players asynchronously choose actions with probability proportional to their exponentiated utility. Due to this desirable property, variants of log-linear learning were proposed in subsequent works (Marden et al., 2007; Marden and Shamma, 2012). For example, Leslie and Marden (2011) proves that log-linear learning also handles the practical setting where the observed utilities are corrupted by noise. Moreover, Arslan et al. (2007) proposes binary log-linear learning, a slight modification of log-linear learning that only requires two points of feedback per round. Furthermore, (Tatarenko, 2017) provides an asymptotic rate of convergence to an efficient NE when the parameter controlling the players’ rationality in log-linear learning is time-varying. Another line of research studies how fast log-linear learning converges to its stationary distribution (Auletta et al., 2010, 2011, 2012) and how it behaves before reaching the stationary distribution (Auletta et al., 2012), but do not focus on Nash equilibria.
Few past works provide finite-time guarantees for log-linear learning to an $\epsilon$-efficient NE, an action profile whose potential is $\epsilon$-close to the maximum value. For instance, for atomic routing games with polynomial costs of degree at most $p$, Asadpour and Saberi (2009) proves a convergence time exponential in $p$ and $1/\epsilon$ and polynomial in $N$, where $N$ denotes the number of players. Moreover, Montanari and Saberi (2008, 2010) study games with graph structures between players and prove an exponential convergence time in $N$ and $1/\epsilon$ in the worst case. Finally, in potential games with interchangeable players and a Lipschitz-continuous potential function, Shah and Shin (2010) shows a convergence time exponential in $A$ and $1/\epsilon$ and linear in $N$, where $A$ is the number of actions per player. The latter result was extended in (Borowski and Marden, 2015) to semi-anonymous potential games, which are potential games with groups of interchangeable players.
Although finite-time convergence of log-linear learning to an $\epsilon$-efficient NE was established for different sub-classes of potential games, such guarantees do not exist for general potential games. In this paper, we focus on deriving the first finite-time convergence guarantees of log-linear learning that hold for general potential games; are polynomial in $1/\epsilon$; and remain valid under relaxed assumptions such as limited access to feedback and noisy utilities. [^2] For general potential games, an exponential dependence on $N$ in the convergence time is unavoidable since finding an $\epsilon$-efficient NE is equivalent to computing an $\epsilon$-optimizer of the potential function, which is NP-complete (Burer and Letchford, 2012).
@@ 1.2 Contributions
  • We prove a convergence time of $\tilde{\mathcal{O}}((A^{N}/\epsilon)^{\frac{1}{\max\{\epsilon,\Delta\}}})$ to an $\epsilon$-efficient NE (Theorem 3.1), where $\Delta$ is a problem-dependent constant. If in addition, the players are interchangeable, then an $\epsilon$-NE is reached in $\tilde{\mathcal{O}}((\frac{N^{A}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$ which in contrast to general potential games is polynomial in $N$ as well (Corollary B.7).
  • We consider two variants, binary log-linear learning (Theorem 4.1) and perturbed log-linear learning (Theorem 4.3) motivated by limited feedback and noise corrupted utilities, respectively. For these variants, we show convergence guarantees that are polynomial in $1/\epsilon$.
On the technical side, past works (Blume, 1993; Young, 1993; Marden and Shamma, 2012; Shah and Shin, 2010) established asymptotic convergence guarantees by leveraging that log-linear learning induces a Markov chain. To obtain our novel finite-time results we build on this connection and develop new Markov chain results that can be summarized as follows:
  • We derive a mixing time bound for a class of Markov chains which includes those induced by log-linear learning and binary-log linear learning. It follows from an adaptation of arguments in (Diaconis and Saloff-Coste, 1996) (Lemma 2.2) and a novel bound on the log-Sobolev constant (Lemma 3.4) that we derive for this class of Markov chains.
  • We derive a tight Lipschitz constant for the known result regarding the Lipschitz-continuity of stationary distributions of Markov chains as a function of their transition matrix (Lemma 4.2). We leverage this result to study the convergence of learning rules such as perturbed log-linear learning for which the explicit stationary distributions are unknown (Theorem 4.3).
@@ 2.1 Problem setup
In this paper, we assume that an initial action profile is drawn from a distribution $\mu^{0}$. The potential game is then repeated over multiple rounds. We assume that at each round $t$, a player $i\in[N]$ is uniformly chosen and selects her action based on some learning rule while the other players remain with their previous action, i.e., $a_{-i}^{t}=a_{-i}^{t-1}$.
A common assumption in game theory is that the players are rational, i.e., that they seek to maximize their utility. Under this assumption, a natural solution concept is the pure Nash equilibrium (Nash Jr, 1950).
Definition 2.1 (Nash equilibrium).
@@
In a pure NE, no player can improve her utility by unilaterally changing her action. Hereafter, we simply refer to a pure Nash equilibrium as a Nash equilibrium.
In potential games, several Nash equilibria may exist as shown in the two-player potential game example below. Here, action profiles $(A,A)$ and $(B,B)$ are both Nash equilibria. However, the value of the potential function may differ for different Nash equilibria, as is the case in the example.
$\bordermatrix{&A&B&\cr A&(5,2)&(-1,-2)\!\!\!\!\!\cr B&(-5,-4)&(1,4)\!\!\!\!\!\!},\quad\text{with the potential:}\quad\bordermatrix{&A&B&\cr A&4&0\!\!\!\!\!\!\!\!\cr B&-6&2\!\!\!\!\!\!\!\!}.$
[…]
In this work, we are interested in learning rules that converge, in expectation, to an $\epsilon$-efficient NE, i.e., an action profile maximizing the potential up to an additive constant:
@@
If all players apply a learning rule that relies exclusively on the utility of the current action profile of the other players then the learning dynamics induce a Markov chain over the state space $\mathcal{A}^{N}$. In particular, the state of the Markov chain at time $t$ corresponds to the action profile at round $t$ and the learning rule of each player specifies the transition dynamics. It follows that we can analyze the convergence time of log-linear learning by studying the associated Markov chain. We, therefore, provide relevant concepts for Markov chains in the next section.
@@ 2.2 Background on Markov chains
We briefly review concepts and properties of Markov chains used throughout this paper. Consider a time-homogeneous Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ over the state space $\mathcal{A}^{N}$ with a transition matrix $P\in\mathbb{R}^{{A^{N}}\times A^{N}}$. The ergodic theorem (Levin and Peres, 2017) states that if a Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ is irreducible and aperiodic, then it has a unique stationary distribution $\mu$, and from any initial distribution $\mu^{0}$ the distribution $\mu^{t}=\mu^{0}P^{t}$ converges to $\mu$. The convergence time to the stationary distribution is quantified by the mixing time:
@@
where the total variation distance is defined as $\|\mu^{t}-\mu\|_{TV}:=\frac{1}{2}\sum_{a\in\mathcal{A}^{N}}|\mu^{t}(a)-\mu(a)|$. Based on a remark in (Diaconis and Saloff-Coste, 1996, Section 3) we now derive a bound on the mixing time of Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$.
Lemma 2.2.
@@
We briefly discuss the mixing time bound above and provide a proof and a thorough discussion in Appendix A. While classical approaches commonly bound the mixing time by the spectral gap defined in Equation (11) in Appendix A, bounds using log-Sobolev constants are often significantly tighter. Indeed, the mixing time upper bound using the log-Sobolev constant grows as $\mathcal{O}(\log\log(1/\mu_{\min}))$ whereas bounds using the spectral gap grow as $\mathcal{O}(\log(1/\sqrt{\mu_{\min}}))$. However, unlike the spectral gap deriving log-Sobolev constants can be extremely difficult. So constants have not been well-explored. In the next section in Lemma 3.4, we derive a novel bound on the log-Sobolev constant of a class of Markov chains.
3 Convergence of log-linear learning
In this section, we review the well-established log-linear learning rule (Blume, 1993) and state our main theoretical result on the convergence time of log-linear learning to an $\epsilon$-efficient Nash equilibrium. Our convergence analysis relies on the mixing time bound from Lemma 2.2 of the previous section.
@@ 3.1 Algorithm and background
[…]
Log-linear learning induces an irreducible and aperiodic Markov chain $\{X_{t}\}_{t\in\mathbb{Z}_{+}}$ with a time-reversible transition matrix $P\in\mathbb{R}^{\mathcal{A}\times\mathcal{A}}$ (Marden and Shamma, 2012) given by:
@@
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. The stationary distribution $\mu\in\Delta(\mathcal{A}^{N})$ of log-linear learning is given by (Blume, 1993):
@@
For sufficiently large $\beta$, it was shown that log-linear learning converges asymptotically to a potential maximizer and thus to an efficient Nash equilibrium (Blume, 1993; Young, 1993; Marden and Shamma, 2012). Except for a few works (Montanari and Saberi, 2010; Asadpour and Saberi, 2009; Shah and Shin, 2010), none of the previous works, however, have finite-time convergence guarantees to such an efficient Nash equilibrium, and (Montanari and Saberi, 2010; Asadpour and Saberi, 2009; Shah and Shin, 2010) have additional assumptions on the potential game. Thus, in the following section, we establish a bound on the convergence time of log-linear learning in general potential games.
3.2 Convergence time
@@
In other words, after running $t=\tilde{\mathcal{O}}(N^{2}A^{5}(\frac{A^{N}}{\epsilon})^{1/\max\{\epsilon,\Delta\}})$ rounds of log-linear learning with $\beta=\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{A^{N}}{\epsilon}$ the expected value of the potential of the joint action at time $t$ is $\epsilon$-optimal.
Theorem 3.1 provides the first finite-time convergence rates to an $\epsilon$-efficient NE in general potential games. In order to be $\epsilon$-close to an efficient NE $\beta$ must scale as $\Omega(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{A^{N}}{\epsilon})$. Furthermore, since the convergence time scales as $e^{4\beta}$ it grows polynomially in $\mathcal{A}$ and $1/\epsilon$ and exponentially in $N$. The exponential dependence on $N$ is unavoidable without further assumptions on the game. However, by deriving problem-dependent bounds we are the first to avoid the exponential dependence in $1/\epsilon$, see Table 1. Note that the case $\Delta=0$ is trivial as all action profiles are efficient NEs. In Appendix E, we conduct numerical experiments to showcase the performance of log-linear learning with $\beta$ specified as in Equation (6). To summarize, our experiments demonstrate that log-linear learning with such a $\beta$ converges to an $\epsilon$-efficient NE. In particular, it can be seen that the suboptimality gap $\Delta$ is an important feature to quantify the convergence time and that the convergence time increases polynomially in $1/\epsilon$ for $\epsilon<\Delta$.
@@
Game setting Assumptions Convergence time
Routing game with $K$ vertices Cost functions of $\tilde{\mathcal{O}}(e^{\frac{N}{\epsilon}})$
(Asadpour and Saberi, 2009) degree at most $p$
Potential game with interchangeable $\lambda$-Lipschitz continuous $\tilde{\mathcal{O}}(N(\frac{A\lambda}{\epsilon})^{\frac{A}{\epsilon}})$
players (Shah and Shin, 2010) potential function
Potential game with interchangeable $A\geq 4$ $\tilde{\mathcal{O}}(N(\frac{N^{A}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$
players (Corollary B.7)
Potential games (Theorem 3.1) $A\geq 4$ $\tilde{\mathcal{O}}(N^{2}A^{5}(\frac{A^{N}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$
Proof 3.2.
outline Here, we provide an outline of the proof, the full proof is deferred to Appendix B.1. The argument begins with the following decomposition:
@@
Second term: We can control the mixing time of log-linear learning using the bound of Lemma 2.2:
@@
Lemma 3.4.
Consider a Markov chain over state space $\mathcal{A}^{N}$ with $A\geq 4$. Assume that there exists $p_{\min},p_{\max}\in(0,1]$, such that the corresponding transition matrix $P\in\mathbb{R}^{A^{N}\times A^{N}}$ satisfies: [^5] Such a $p_{\min}\in(0,1]$ exists if the probability of transitioning from any state to a neighboring state is positive.
@@
for $t\geq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\left(\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\right)$. This concludes the proof of the theorem.
In the following, we additionally assume that the potential game is symmetric, i.e., players are interchangeable. [^6] A definition of a symmetric game is given in Appendix B.4, see Definition B.6. In these games, a player’s utility depends solely on the number of players selecting each action and not on their identity. Technically speaking, in this case, log-linear learning induces a Markov chain on a modified state space of cardinality $\mathcal{O}(N^{A})$, see Equation (22), rather than on the state space $A^{N}$. Exploiting the smaller state space, we derive a bound on the convergence time of log-linear learning that depends polynomially not only on $1/\epsilon$ but also on $N$.
Corollary 3.5 (Sketch).
Consider a symmetric potential game. For any $\epsilon\in(0,1)$ and initial distribution $\mu^{0}$, assume that players adhere to modified log-linear learning (Shah and Shin, 2010) with $\beta=\Omega(\frac{1}{\max\{\epsilon,\Delta\}}\log(\frac{N^{A}}{\epsilon}))$. Then, $\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon$ for $t=\tilde{\mathcal{O}}(N(\frac{N^{A}}{\epsilon})^{\frac{1}{\max\{\epsilon,\Delta\}}})$.
We provide a full statement of the corollary and its proof in Appendix B.4. For symmetric potential games with a $\lambda$-Lipschitz-continuous potential function, Shah and Shin (2010) proves a convergence time of $\tilde{\mathcal{O}}(N(\frac{A\lambda}{\epsilon})^{\frac{A}{\epsilon}})$. In comparison, relaxing the Lipschitzness assumption comes at a small cost of polynomial dependence on $N$ rather than a linear one. However, our result greatly improves the dependence on $\epsilon$ to a polynomial one.
4 Robustness of log-linear learning
In this section, we show that our convergence time guarantee is robust to settings where players have less feedback information and to settings where log-linear learning is subject to perturbations such as noisy utility observations.